OpenStudy (anonymous):

I toss a coin 4 times. Find the chance of getting 2 heads.

OpenStudy (raden):

it means the probability of P(2H and 2T) first, find the number events of HHTT, are 4!/(2!2!) = 6 we knowed that the probability of head event = 1/2 also tail even = 1/2 and it is independent, so, it can be P(2H and 2T) = 6 * 1/2 * 1/2 * 1/2 * 1/2 = 6/16 = 3/8